# Dynamic planning in hierarchical active inference

### v2

@@ Abstract
By dynamic planning, we refer to the ability of the human brain to infer and impose motor trajectories related to cognitive decisions. A recent paradigm, active inference, brings fundamental insights into the adaptation of biological organisms, constantly striving to minimize prediction errors to restrict themselves to life-compatible states. Over the past years, many studies have shown how human and animal behaviors could be explained in terms of an active inferential process – either as discrete decision-making or continuous motor control – inspiring innovative solutions in robotics and artificial intelligence. Still, the literature lacks a comprehensive outlook on how to effectively plan realistic actions in changing environments. Setting ourselves the goal of modeling tool use, we delve into the topic of dynamic planning in active inference, keeping in mind two crucial aspects of biological goal-directed behavior: the capacity to understand and exploit affordances for object manipulation, and to learn the hierarchical interactions between the self and the environment, including other agents. We start from a simple unit and gradually describe more advanced structures, comparing recently proposed design choices and providing basic examples for each section. This study distances itself from traditional views centered on neural networks and reinforcement learning, and points toward a yet unexplored direction in active inference: hybrid representations in hierarchical models.
@@ 1 Introduction
Hierarchies are found everywhere in the world. They are so pervasive that they do not just exist as causal relationships between physical properties of the environment, but are also inherent to how biological organisms act over it. Even the most complex kinematic structures of animals follow a rigid hierarchical strategy, whereby different limbs propagate from a body-centered reference frame. The hierarchical modularity of brain functional networks is widely recognized [1, 2], as well as the representation of the body schema in somatosensory and motor areas [3], and the organization of hierarchical motor sequences concerning parietal and premotor cortices [4]. In particular, the body schema is not a static entity but changes in concurrence to the development of the human body during childhood and adolescence [5]. Surprisingly, the nervous system is able to relate external objects to the self in a way that, although not reflecting the actual causal relationships between the body and the environment, is the most suitable for better operating in a specific context. Physiological studies have demonstrated that, with extensive tool use, parietal and motor areas of the monkey brain gradually adapt to make room for the tool, increasing the length of the perceived limb [6, 7]. This adaptation is highly plastic, assimilating objects in a very short time [8] and inducing altered somatosensory representations of the body morphology that persist even after tool use [9].
Why and how does this happen? One recent theory is that of predictive coding [10, 11, 12], which has been attracting increasing interest in recent years and proposes itself as a unifying paradigm of cortical function. Predictive coding posits that living beings make sense of the world by building an internal generative model that tries to imitate the hierarchical causal relationships of the external generative process. From a high-level hypothesis about the state of affairs of the world, a cascade of neural predictions takes place, eventually leading to a low-level guess about sensory evidence. Comparing the model’s guess with the sensorium triggers another cascade of prediction errors that travel back to the deepest cortical levels. The model iteratively refines its structure until all the prediction errors are minimized, that is, until it is finally able to predict what will happen next. This optimization differs from the more traditional view of deep learning, in that the message passing is local and what climbs up the hierarchy does not signal the detection of a feature, but how much the model is surprised about its prediction. Besides having stimulated cognitive and neural studies under several circumstances [13, 14, 15, 16], this theory has also influenced novel directions in machine learning: Predictive Coding Networks (PCNs) have been shown to generalize well to classification or regression tasks [17, 18], with key advantages compared to neural networks and still approximating the backpropagation algorithm [19, 20, 21].
@@
While predictive coding can elucidate – through a rigorous computational framework [22] – illusions and visual phenomena such as binocular rivalry [23], it explains just the first (perceptual) half of the story. More specifically, it does not explain why interactions with the environment occur – a process that results, considering the above example, in the monkey brain actively distorting its body schema during tool use. Such complex tasks always involve decision-making, which the brain is known to realize via several methods [24]. Among them, one is particularly relevant here: planning for deliberation, also known as vicarious trials and errors, whereby an action is selected after several alternatives have been generated and evaluated [25]. One of the most intriguing characteristics of human planning is the capacity to imagine, or endogenously generate dynamic representations of future states, including potential trajectories and subgoals that bring to such states [26, 27]. The hippocampus is a key neural structure known to support trajectory generation, although planning is accomplished in concert with other areas implementing evaluation of options and response selection [25]. How does the human brain account for the dynamics of the self and the environment to afford purposeful planning?
@@
On this trail, a second innovative perspective has been proposed, aspiring to unveil a unified first principle not just on cortical function, but on the behavior of all living organisms. This perspective, called active inference [28, 29, 30, 31], is grounded on the same theoretical basis of predictive coding but further assumes two key aspects of biological behavior. First, that a living being does not maintain static hypotheses about the state of affairs of the world but can also construct internal dynamics – either as instantaneous trajectories or future states – allowing it to anticipate the unfolding of events occurring at different timescales. Second, that these dynamic hypotheses can be fulfilled by movements. The latter assumption replaces models with agents, conveying a somewhat counterintuitive but insightful implication: while perception lets the agent’s hypothesis conform to the environment (as in predictive coding), action forces the environment to conform to the hypothesis, by letting the agent sample those observations that make the hypothesis true. If such hypotheses (usually called beliefs) correspond to desired states defined, e.g., by the phenotype, cycling between action and perception ultimately allows the agent to survive. This is the core of the so-called free energy principle, which states that in order to maintain homeostasis, all organisms must constantly and actively minimize the difference between their sensory states and their expectations based on a small set of life-compatible options. Giving a practical example, if I believe to find myself with a tool in hand, I will try with all my strength to observe visual images of the tool in my hand; in doing this, a combined reaching and grasping action happens. This view distances itself from the stimulus-response mapping widely estabilished in neuroscience, and evidence indicates that it could be more biologically plausible than optimal control and Reinforcement Learning (RL) [32, 33, 34, 35].
In principle, active inference might be key for understanding how goal-directed behavior emerges in the human brain [36]. For instance, relevant objects used for manipulation may gradually become part of one’s identity through a closed loop between motor commands and sensory evidence, meaning that the boundary of the self from the environment increases whenever the agent manages to predict the consequences of its own movements [37]. Additionally, active inference might prove fundamental for making advances with current artificial agents, taking forward a promising research area known as planning as inference [38, 39, 40]. Active inference implementations can be divided into two frameworks, which have been used to simulate human and animal behaviors under the two complementary aspects of motor control [32, 41, 42, 43, 44, 45] and decision-making [46, 47, 48, 49, 50]. The first framework – generally compared to the low-level sensorimotor loops – is defined in continuous time [51, 52] and makes use of generalized filtering [53] to model instantaneous trajectories of the self and the environment; these trajectories are inferred by minimization of a quantity called variational free energy, which is the negative of what in machine learning is known as the evidence lower bound. Differently from optimal control, motor commands in active inference derive from proprioceptive predictions that are fulfilled by classical spinal reflex arcs [34]. This eliminates the need for cost functions – as the inverse model maps from proprioceptive (and not latent) states to actions – and replaces a control problem with an inference problem [33]. The second framework – attributed to the cerebral cortex, especially prefrontal aras [54], along with corticostriatal loops – is expressed in discrete state-space [55, 56] and exploits the structure of Partially Observable Markov Decision Processes (POMDPs) to plan abstract actions over expected future sensations. This (active) inference relies on the minimization of the expected free energy, i.e., the free energy that the agent expects to perceive in the future. The expected free energy can be unpacked into two terms resembling the two classical aspects of control theory, exploration and exploitation – which here naturally arise; these respectively correspond to an uncertainty-reducing term, and a goal-seeking term that, as before, pushes the agent to find a sequence of actions leading to its prior belief.
Three features of active inference are relevant to designing intelligent agents that can tackle real-life applications and, for the goal of this study, tasks requiring tool use. First, multiple units – composed of simple likelihood and state transition distributions – can be easily connected to adapt to complex hierarchical structures [57]. For instance, a hierarchical kinematic model can be designed in continuous time, wherein each unit encodes a certain Degree of Freedom (DoF) in intrinsic and extrinsic reference frames [44]. This allows one to realize advanced movements that involve simultaneous coordination of multiple limbs, e.g., moving with a glass in hand. This hierarchical structure can be generalized to perform homogeneous transformations between reference frames, e.g., perspective projections [58]. However, a continuous model alone lacks effective usability in the real world, since it can only deal with present sensory states and cannot perform any form of future planning.
The latter is usually possible through so-called mixed or hybrid models [48, 59], which combine the potentialities of a discrete model with the inference of continuous signals, allowing robust decision-making in uncertain and changing environments. While the theory of Bayesian model reduction [60, 61, 62, 63] provides efficient communication between the two models, this unified approach has not enjoyed many practical implementations for the time being [31, 48, 59, 64, 65, 66, 67]. An open issue regards how to deal with highly dynamic environments: standard hybrid models usually perform a comparison between static priors, limiting the agent to realize, e.g., multi-step reaching movements through fixed positions. One study addressed the problem of realistic robot navigation in active inference, but making use of alternative bio-inspired SLAM methods [68]. In [69], a hybrid model in which the agent’s hypotheses were generated at each time step from the system dynamics allowed to relate continuous trajectories with discrete plans.
A third appealing characteristic of the framework is that one can encode beliefs not only over its own bodily states, but also over external physical variables. This has been recently done in the context of active object reconstruction [70, 71, 72, 73] – where an agent encoded independent representations for multiple elements, and used action to more accurately infer its dynamics; for simulating oculomotor behavior [74] – where the dynamics of a target belief was biased by a hidden location; or for analyzing epistemic affordance [50], i.e., the changes in affordance of different objects in relation to the agent’s beliefs. In continuous time, such affordances can be expressed in intrinsic reference frames corresponding to potential agent’s configurations, defining specific ways to interact with the objects. Manipulating these additional beliefs depending on the agent’s intentions [43] permits effectively operating in dynamic contexts, e.g., tracking a target with the eyes [74], or grasping an object on the fly [69] and placing it at a goal position [75]. However, these applications do not exploit the efficiency of deep hierarchical models, and controlling multiple limbs other than the hand is not straightforward. Crucially, they would not appreciate the flexibility of animal brains in remapping their neuronal activity to account for usable tools.
Based on these premises, a question arises: how to perform dynamic planning with hierarchical structures of several objects? In other words, how to combine these three features into a single view? While many studies in continuous time can be currently found in the literature [52, 76, 77], a rigorous formalism of how to realize goal-directed behavior is still lacking, with the consequence of using different solutions for similar problems – especially in contexts that demand online replanning. On the other side, promising results have been achieved by combining the capabilities of discrete-time active inference with neural networks, in a way loosely resembling deep RL. Indeed, so-called deep active inference have critical advantages in learning and solving online tasks compared to traditional methods [78, 79, 80, 81, 82]. Still, neural networks are treated as black boxes during free energy minimization, without fully enjoying the potential benefits of hierarchical and temporal depths. One of the most attractive aspects of active inference is that it prescribes a unified perspective not just for fitting to complex high-dimensional data (as neural networks or PCNs do), but also for embodying environmental dynamics and acting over them to minimize uncertainty and conform to prior beliefs.
@@
For these reasons, in this study we explore an alternative direction to the optimal control problem that does not call for additional frameworks, in a few words, a direction toward hybrid computations in hierarchical systems. We analyze many design choices that have been applied in the motor control domain, with an in-depth look at the three characteristics mentioned above. Asking ourselves how to model tool use, we start from a simple unit and construct richer modules that can be linked in a hierarchical fashion, exhibiting interesting high-level features. In Chapter 2, we consider a single-DoF agent and explore how to realize a sort of multi-step behavior in continuous time only. In Chapter 3, we analyze the implications of combining different units in a single network, using more complex kinematic configurations and distinguishing between intrinsic and extrinsic dynamics. In Chapter 4, we describe the advantages of using discrete decision-making in continuous environments, focusing on hybrid structures and drawing some parallelisms between the two worlds. Finally, in the Discussion we elaborate on the benefits of addressing discrete and continuous representations together, and give a few suggestions for future work on this subject.
2 Flexible intentions
@@
In this chapter, we begin by explaining the inferential mechanisms of a basic unit in continuous time. We then discuss one by one the changes and features that we introduce, in order to achieve a multi-step behavior in simple tasks that do not require deep hierarchical modeling nor online replanning.
@@ 2.1 A simple agent
Figure 1: Factor graph of a basic unit for static reaching. Variables and factors are indicated by circles and squares, respectively. Hidden states $\bm{x}$ (e.g., the hand position) generate observations $\bm{o}$ (e.g., an image of the hand) through the likelihood function $\bm{g}$, and their 1st derivatives $\bm{x}^{\prime}$ (e.g., the hand velocity) through a dynamics function $\bm{f}$. In contrast to optimal control, here action follows observation prediction errors arising from a simple attractor $\bm{\rho}$ embedded in the model dynamics, or from a prior belief $\bm{\eta}$ over the hand position.
The most elementary unit is represented in Figure 1. This is the simplest formulation of a continuous-time active inference agent, where we kept only the key nodes. This allows us to easily describe a velocity-controlled dynamic system with the following likelihood $\bm{g}$ and dynamics $\bm{f}$:
$\begin{split}\bm{o}&=\bm{g}(\bm{x})+\bm{w}_{o}\\ \bm{x}^{\prime}&=\bm{f}(\bm{x})+\bm{w}_{x}\end{split}$ (1)
where $\bm{x}$ and $\bm{o}$ are respectively called hidden states and observations, and the letter $w$ indicates noise terms sampled from Gaussian distributions. For simplicity, we considered just two temporal orders – although all the features that we elucidate in the following hold for a system of generalized coordinates [53] – and we defined a likelihood function only for a single temporal order. We assume that the corresponding generative model is factorized as follows:
$p(\tilde{\bm{x}},\bm{o})=p(\bm{o}|\bm{x})p(\bm{x})p(\bm{x}^{\prime}|\bm{x})$ (2)
where:
@@
[…]
$D_{KL}[q(\tilde{\bm{x}})||p(\tilde{\bm{x}}|\bm{o})]=\int_{\tilde{\bm{x}}}q(\tilde{\bm{x}})\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}}|\bm{o})}d\tilde{\bm{x}}$ (4)
@@
The denominator $p(\bm{x}|\bm{y})$ still depends on the marginal $p(\bm{y})$, but the KL divergence can be rewritten in terms of the log evidence and a quantity known as the free energy $\mathcal{F}$:
$\mathcal{F}=\operatorname*{\mathbb{E}}_{q(\tilde{\bm{x}})}\left[\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}},\bm{o})}\right]=\operatorname*{\mathbb{E}}_{q(\tilde{\bm{x}})}\left[\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}}|\bm{o})}\right]-\ln p(\bm{o})$ (5)
@@
Since the KL divergence is always nonnegative, the free energy provides an upper bound on surprise, i.e., $\mathcal{F}\geq\ln p(\bm{o})$. Hence, minimizing the KL divergence with respect to $q(\tilde{\bm{x}})$ is equivalent to minimizing $\mathcal{F}$, and achieves the dual objective of keeping surprise low while estimating the true distribution. Assuming that the approximate posterior can be factorized into independent contributions, and further assuming that each contribution is Gaussian, the optimization process breaks down to the minimization of the prediction errors associated with the distributions of the generative model in Equation 2 – see [31] for more details:
$\begin{split}\bm{\varepsilon}_{o}&=\bm{o}-\bm{g}(\bm{\mu})\\ \bm{\varepsilon}_{\eta}&=\bm{\mu}-\bm{\eta}\\ \bm{\varepsilon}_{x}&=\bm{\mu}^{\prime}-\bm{f}(\bm{\mu})\end{split}$ (6)
@@
Then, the inference of the means (also called beliefs) of the posterior over the hidden states, denoted by $\tilde{\bm{\mu}}$, is reduced to the following message passing:
$\dot{\tilde{\bm{\mu}}}=\begin{bmatrix}\dot{\bm{\mu}}\\ \dot{\bm{\mu}}^{\prime}\end{bmatrix}=\mathcal{D}\tilde{\bm{\mu}}-\partial_{\tilde{\mu}}\mathcal{F}=\begin{bmatrix}\bm{\mu}^{\prime}-\bm{\pi}_{\eta}\bm{\varepsilon}_{\eta}+\partial_{\mu}\bm{g}^{T}\bm{\pi}_{o}\bm{\varepsilon}_{o}+\partial_{\mu}\bm{f}^{T}\bm{\pi}_{x}\bm{\varepsilon}_{x}\\ \\ -\bm{\pi}_{x}\bm{\varepsilon}_{x}\end{bmatrix}$ (7)
@@
where $\mathcal{D}$ is an operator that shifts every derivative by one, i.e., $\mathcal{D}\tilde{\bm{\mu}}=[\bm{\mu}^{\prime},\bm{0}]$. This term arises because the generative model maintains a belief not over a static point, but over a dynamic trajectory, and only when the motion of the mean $\dot{\tilde{\bm{\mu}}}$ equals the mean of the motion $\mathcal{D}\tilde{\bm{\mu}}$, is the free energy minimized. In short, the inferential process does not involve matching a state (as in PCNs) but tracking a path [84]. Unpacking Equation 7, we note that the 0th order is subject to a forward error from the prior, a backward error from the likelihood, and a backward error from the dynamics function. On the other hand, the 1st order is only subject to the latter but in the form of a forward error. The belief is then updated via gradient descent, i.e., $\tilde{\bm{\mu}}_{t+1}=\tilde{\bm{\mu}}_{t}+\Delta_{t}\dot{\tilde{\bm{\mu}}}$, where $\Delta_{t}$ is a time constant.
@@
How can this agent perform a simple reaching movement? As highlighted in Figure 1, we can encode the hand position and velocity as generalized hidden states. We will talk later about the relation between proprioceptive and exteroceptive domains, as this deserves a careful discussion. For now, we consider a single DoF that has a univocal mapping between the joint angle and the Cartesian position, so we can represent both of them by the same variables and factors (note however that we maintain a bold notation for generalization and consistency with the rest of the study). Indicating the target to reach by $\bm{\rho}$, we can define the following dynamics function:
@@
expressing a simple attractor toward the target [37, 85, 86, 87, 88]. This dynamics does not exist in the actual generative process, and it is indeed this discrepancy that forces the environment to conform to the agent’s beliefs. Specifically, Equation 8 means that the agent thinks its hand will be pulled toward the target with a strength proportional to the precision $\bm{\pi}_{x}$. In fact, the attractor affects the belief update through the dynamics prediction error $\bm{\varepsilon}_{x}$, expressing a difference between the estimated velocity $\bm{\mu}^{\prime}$ and the one predicted by the agent through the dynamics function $\bm{f}$. Note that this error appears in both temporal orders: in brief, $\bm{\varepsilon}_{x}$ imposes a trajectory at the 1st order which in turn affects the 0th order directly through $\bm{\mu}^{\prime}$, and indirectly through the gradient $\partial_{\mu}\bm{f}$.

### v3

@@ Abstract
By dynamic planning, we refer to the ability of the human brain to infer and impose motor trajectories related to cognitive decisions. A recent paradigm, active inference, brings fundamental insights into the adaptation of biological organisms, constantly striving to minimize prediction errors to restrict themselves to life-compatible states. Over the past years, many studies have shown how human and animal behaviors could be explained in terms of active inference – either as discrete decision-making or continuous motor control – inspiring innovative solutions in robotics and artificial intelligence. Still, the literature lacks a comprehensive outlook on effectively planning realistic actions in changing environments. Setting ourselves the goal of modeling complex tasks such as tool use, we delve into the topic of dynamic planning in active inference, keeping in mind two crucial aspects of biological behavior: the capacity to understand and exploit affordances for object manipulation, and to learn the hierarchical interactions between the self and the environment, including other agents. We start from a simple unit and gradually describe more advanced structures, comparing recently proposed design choices and providing basic examples. This study distances itself from traditional views centered on neural networks and reinforcement learning, and points toward a yet unexplored direction in active inference: hybrid representations in hierarchical models.
@@ 1 Introduction
Three characteristics of the brain are relevant to tasks involving planning in changing environments, such as tool use. First, the brain’s ability to maintain estimates not only of bodily states, but also of external physical variables in relation to the self. Studies have shown how the Posterior Parietal Cortex (PPC) of the monkey brain encodes objects with sensorimotor representations reflecting the body structure [1, 2]. These representations are extremely useful for object manipulation since they account efficiently for the action possibilities provided by the object, also known as affordances [3]. For instance, one might encode a cup in different ways (power versus precision grips) based on whether one wants to throw it or drink from it. Further, to act timely in a dynamic environment, the PPC can encode multiple objects in parallel during sequences of actions, even when there is a considerable delay between different subgoals [4].
@@
A second characteristic regards flexible and deep hierarchies. Hierarchical structures are so pervasive that they not only exist as causal relationships between physical properties of the environment, but are also inherent to how biological organisms act over it. Even the most complex kinematic structures of animals follow a rigid hierarchical strategy, whereby different limbs propagate from a body-centered reference frame. The hierarchical modularity of brain functional networks is widely recognized [5, 6], as well as the representation of the body schema in somatosensory and motor areas [7], and the organization of hierarchical motor sequences concerning parietal and premotor cortices [8]. In particular, the body schema is not a static entity but changes in concurrence to the development of the human body during childhood and adolescence [9]. Surprisingly, the nervous system is able to relate external objects to the self in a way that, although not reflecting the actual causal relationships between the body and the environment, is the most suitable for better operating in a specific context. Physiological studies have demonstrated that, with extensive tool use, parietal and motor areas of the monkey brain gradually adapt to make room for the tool, increasing the length of the perceived limb [10, 11]. This adaptation is highly plastic, assimilating objects in a very short time [12] and inducing altered somatosensory representations of the body morphology that persist even after tool use [13].
@@
A third characteristic is the ability to construct a dynamic discretized plan based on continuous sensory evidence. Complex tasks involve decision-making, which the brain is known to realize via several methods [14]. Among them, one is particularly relevant: planning for deliberation, also known as vicarious trials and errors, whereby an action is selected after several alternatives have been generated and evaluated [15]. One of the most intriguing aspects of human planning is the capacity to imagine, or endogenously generate dynamic representations of future states, including potential trajectories and subgoals that bring to such states [16, 17]. The hippocampus is a key neural structure known to support trajectory generation, although planning is accomplished in concert with other areas [15].
How does the human brain capture the hierarchical organization and dynamics of the self and the environment to afford purposeful planning? One recent theory is that of predictive coding [18, 19, 20, 21], which has been attracting increasing interest in recent years and proposes itself as a unifying paradigm of cortical function. According to predictive coding, living beings make sense of the world by building an internal generative model that imitates the causal relationships of the external generative process. From a high-level hypothesis about the world, a cascade of neural predictions takes place, eventually leading to a low-level guess about sensory evidence. Comparing the model’s guess with the sensorium triggers another cascade of prediction errors that travel back to the deepest cortical levels. The model iteratively refines its structure until all the prediction errors are minimized, that is, until it correctly predicts what will happen next. This optimization differs from the more traditional view of deep learning, in that the message passing is local and what climbs up the hierarchy does not signal the detection of a feature, but how much the model is surprised about its prediction. Besides having stimulated cognitive and neural studies under several circumstances [22, 23, 24, 25], this theory has also influenced novel directions in machine learning: Predictive Coding Networks (PCNs) have been shown to generalize well to classification or regression tasks [26, 27], with key advantages compared to neural networks and still approximating the backpropagation algorithm [28, 29, 30].
@@
While predictive coding can elucidate illusions and visual phenomena such as binocular rivalry [31], it explains just the first (perceptual) half of the story. More specifically, it does not explain why interactions with the environment occur – a process that results, considering the above example, in the monkey brain actively distorting its body schema during tool use. On this trail, a second innovative perspective has been proposed, aspiring to unveil a unified first principle not just on cortical function, but on the behavior of all living organisms. This perspective, called active inference [32, 33, 34, 35], is grounded on the same theoretical basis of predictive coding but further assumes two key aspects of biological behavior. First, that a living being does not maintain static hypotheses about the world but also constructs internal dynamics – either as instantaneous trajectories or future states – to anticipate the unfolding of events occurring at different timescales. Second, that these dynamic hypotheses can be fulfilled by movements. The latter assumption replaces models with agents, conveying a somewhat counterintuitive but insightful implication: while perception lets the agent’s hypothesis conform to the environment (as in predictive coding), action forces the environment to conform to the hypothesis – by sampling those observations that make the hypothesis true. If such hypotheses or beliefs correspond to desired states defined, e.g., by the phenotype, cycling between action and perception ultimately allows the agent to survive. This is the core of the so-called free energy principle, which states that in order to maintain homeostasis, all organisms must constantly and actively minimize the difference between their sensory states and their expectations based on a small set of life-compatible options. Giving a practical example, if I believe to find myself with a tool in hand, I will try with all my strength to observe visual images of the tool in my hand; in doing this, a combined reaching and grasping action happens. This view distances itself from the stimulus-response mapping widely established in neuroscience, and evidence indicates that it could be more biologically plausible than optimal control and Reinforcement Learning (RL) [36, 37, 38, 39].
Active inference implementations can be divided into two frameworks, which have been used to simulate human and animal behaviors under the two complementary aspects of motor control [40, 41, 42, 43, 44, 45, 46] and decision-making [47, 48, 49, 50, 51]. In principle, active inference might be key for understanding how goal-directed behavior emerges in the human brain [52]. For instance, relevant objects used for manipulation may gradually become part of one’s identity through a closed loop between motor commands and sensory evidence, meaning that the boundary of the self from the environment increases whenever the agent predicts the consequences of its own movements [53]. Additionally, active inference might lead to key advances with current artificial agents, taking forward a promising research area known as planning as inference [54, 55, 56]. The three characteristics delineated above are fundamental to designing active inference agents that can tackle real-life applications such as tool use. But how to combine them into a single view? In other words, how to perform dynamic planning with hierarchical structures of several objects?
@@
To answer this question, in this study we explore an alternative direction in active inference, i.e., toward hybrid computations in hierarchical systems. We analyze many design choices that have been applied in the motor control domain, with an in-depth look at object affordances, deep hierarchies, and planning with continuous signals. Asking ourselves how to model tool use, we start from a simple unit and construct richer modules that can be linked in a hierarchical fashion, exhibiting interesting high-level features. In Chapter 2, we consider a single-DoF agent and explore how to account for affordances and realize a multi-step behavior in continuous time only. In Chapter 3, we analyze the implications of combining different units in a single network, using more complex kinematic configurations and distinguishing between intrinsic and extrinsic dynamics. In Chapter 4, we describe the advantages of using discrete decision-making in continuous environments, focusing on hybrid structures and drawing some parallelisms between the two worlds. Finally, in the Discussion we elaborate on the benefits of addressing discrete and continuous representations together, and give a few suggestions for future work on this subject.
2 Modeling affordances
@@
In this chapter, we explain the inference mechanisms of a basic unit in continuous time. We then discuss one by one the changes and features that we introduce, in order to achieve a multi-step behavior in simple tasks that do not require deep hierarchical modeling nor online replanning.
The continuous-time active inference framework [57, 36, 40] – generally compared to the low-level sensorimotor loops – makes use of generalized filtering [58] to model instantaneous trajectories of the self and the environment; these trajectories are inferred by minimization of a quantity called variational free energy, which is the negative of what in machine learning is known as the evidence lower bound. Differently from optimal control, motor commands in active inference derive from proprioceptive predictions that are fulfilled by classical spinal reflex arcs [38]. This eliminates the need for cost functions – as the inverse model maps from proprioceptive (and not latent) states to actions – and replaces a control problem with an inference problem [37].
Modeling of objects in active inference has been recently done in the context of active object reconstruction [59, 60, 61, 62] – where an agent encoded independent representations for multiple elements, and used action to more accurately infer its dynamics; for simulating oculomotor behavior [63] – where the dynamics of a target belief was biased by a hidden location; or for analyzing epistemic affordance [51], i.e., the changes in affordance of different objects in relation to the agent’s beliefs. In continuous time, such affordances can be expressed in intrinsic reference frames corresponding to potential agent’s configurations, defining specific ways to interact with the objects. Manipulating these additional beliefs depending on the agent’s intentions [64] permits effectively operating in dynamic contexts, e.g., tracking a target with the eyes [63], or grasping an object on the fly [65] and placing it at a goal position [66].
@@ 2.1 A simple agent
Figure 1: (a) Factor graph of a basic unit for static reaching. Variables and factors are indicated by circles and squares, respectively. Hidden states $\bm{x}$ (e.g., the arm angle) generate observations $\bm{o}_{p}$ (e.g., the arm proprioception) through the likelihood function $\bm{g}_{p}$, and their 1st derivatives $\bm{x}^{\prime}$ (e.g., the arm velocity) through a dynamics function $\bm{f}$. In contrast to optimal control, here action follows observation prediction errors arising from a simple attractor $\bm{\rho}$ embedded in the model dynamics, or from a prior belief $\bm{\eta}_{x}$ over the arm angle. (b) Agent’s generative model.
The most elementary unit is represented in Figure 1(a). This is the simplest formulation of a continuous-time active inference agent, where we kept only the key nodes. This allows us to easily describe a velocity-controlled dynamic system with the following likelihood $\bm{g}_{p}$ and dynamics $\bm{f}$:
$\begin{split}\bm{o}_{p}&=\bm{g}_{p}(\bm{x})+\bm{w}_{o,p}\\ \bm{x}^{\prime}&=\bm{f}(\bm{x})+\bm{w}_{x}\end{split}$ (1)
@@
$D_{KL}[q(\tilde{\bm{x}})||p(\tilde{\bm{x}}|\bm{o}_{p})]=\int_{\tilde{\bm{x}}}q(\tilde{\bm{x}})\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}}|\bm{o}_{p})}d\tilde{\bm{x}}$ (2)
@@
The denominator $p(\bm{x}|\bm{o}_{p})$ still depends on the marginal $p(\bm{o}_{p})$, but the KL divergence can be rewritten in terms of the log evidence and a quantity known as the free energy $\mathcal{F}$:
$\mathcal{F}=\operatorname*{\mathbb{E}}_{q(\tilde{\bm{x}})}\left[\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}},\bm{o}_{p})}\right]=\operatorname*{\mathbb{E}}_{q(\tilde{\bm{x}})}\left[\ln\frac{q(\tilde{\bm{x}})}{p(\tilde{\bm{x}}|\bm{o}_{p})}\right]-\ln p(\bm{o}_{p})$ (3)
@@
Since the KL divergence is always nonnegative, the free energy provides an upper bound on surprise, i.e., $\mathcal{F}\geq\ln p(\bm{o}_{p})$. Hence, minimizing $\mathcal{F}$ achieves the dual objective of keeping surprise low while estimating the true distribution. Assuming that the approximate posterior can be factorized into independent contributions, and further assuming that each contribution is Gaussian – i.e., $\bm{q}(\tilde{\bm{x}})=\mathcal{N}(\tilde{\bm{\mu}}_{x},\tilde{\bm{P}}_{x})$, with generalized means $\tilde{\bm{\mu}}_{x}$ and precisions $\tilde{\bm{P}}_{x}$ – the optimization process breaks down to the minimization of (precision-weighted) prediction errors in terms of the approximate posterior – see [35] for more details:
$\begin{split}\bm{\varepsilon}_{o,p}&=\bm{o}_{p}-\bm{g}_{p}(\bm{\mu}_{x})\\ \bm{\varepsilon}_{\eta,x}&=\bm{\mu}_{x}-\bm{\eta}_{x}\\ \bm{\varepsilon}_{x}&=\bm{\mu}_{x}^{\prime}-\bm{f}(\bm{\mu}_{x})\end{split}$ (4)
@@
Then, the inference of the means $\tilde{\bm{\mu}}_{x}=[\bm{\mu}_{x},\bm{\mu}_{x}^{\prime}]$ (also called beliefs) of the posterior over the hidden states is reduced to the following message passing:
$\dot{\tilde{\bm{\mu}}}_{x}=\begin{bmatrix}\dot{\bm{\mu}}_{x}\\ \dot{\bm{\mu}}^{\prime}_{x}\end{bmatrix}=\mathcal{D}\tilde{\bm{\mu}}_{x}-\partial_{\tilde{x}}\mathcal{F}=\begin{bmatrix}\bm{\mu}_{x}^{\prime}-\bm{\Pi}_{\eta,x}\bm{\varepsilon}_{\eta,x}+\partial_{x}\bm{g}_{p}^{T}\bm{\Pi}_{o,p}\bm{\varepsilon}_{o,p}+\partial_{x}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}\\ \\ -\bm{\Pi}_{x}\bm{\varepsilon}_{x}\end{bmatrix}$ (5)
@@
where $\mathcal{D}$ is an operator that shifts every derivative by one, i.e., $\mathcal{D}\tilde{\bm{\mu}}_{x}=[\bm{\mu}_{x}^{\prime},\bm{0}]$. This term arises because the generative model maintains a belief not over a static point, but over a dynamic trajectory, and only when the motion of the mean $\dot{\tilde{\bm{\mu}}}_{x}$ equals the mean of the motion $\mathcal{D}\tilde{\bm{\mu}}_{x}$, is the free energy minimized. In short, the inferential process does not involve matching a state (as in PCNs) but tracking a path [69]. Unpacking Equation 5, we note that the 0th order is subject to a forward error from the prior, a backward error from the likelihood, and a backward error from the dynamics function. On the other hand, the 1st order is only subject to the latter but in the form of a forward error. The belief is then updated via gradient descent, i.e., $\tilde{\bm{\mu}}_{x,t+1}=\tilde{\bm{\mu}}_{x,t}+\Delta_{t}\dot{\tilde{\bm{\mu}}}_{x}$, where $\Delta_{t}$ is a time constant.
@@
How can this agent perform a simple reaching movement? As highlighted in Figure 1(a), we can encode the arm angle and velocity as generalized hidden states. We will talk later about the relation between proprioceptive and exteroceptive domains; for now, we consider a single DoF that has a univocal mapping between the joint angle of the arm and the Cartesian position of the hand. Indicating the target to reach by $\bm{\rho}$, we can define the following dynamics function:
@@
expressing a simple attractor toward the target [53, 70, 71, 72, 73]. These dynamics do not exist in the actual generative process, and it is indeed this discrepancy that forces the environment to conform to the agent’s beliefs. Specifically, Equation 6 means that the agent thinks its hand will be pulled toward the target with a strength proportional to the precision $\bm{\Pi}_{x}$. In fact, the attractor affects the belief update through the dynamics prediction error $\bm{\varepsilon}_{x}$, expressing a difference between the estimated velocity $\bm{\mu}_{x}^{\prime}$ and the one predicted by the agent through the dynamics function $\bm{f}$. Note that this error appears in both temporal orders: in brief, $\bm{\varepsilon}_{x}$ imposes a trajectory at the 1st order which in turn affects the 0th order directly through $\bm{\mu}_{x}^{\prime}$, and indirectly through the gradient $\partial_{x}\bm{f}$.